An engine does 3250 J of work and discards 5750 J of heat. Find the efficiency of this engine.


1
Expert's answer
2021-10-31T18:09:38-0400

Let's first find the quantity of heat that supplied to the engine in each cycle:


"W=Q_H-Q_C,""Q_H=W+Q_C=3250\ J+5750\ J=9000\ J."

Finally, we can find the efficiency of the engine:


"e=\dfrac{W}{Q_H}\times100\%=\dfrac{3250\ J}{9000\ J}\times100\%=36\%."